If the area of a rectangle is 50 inches and lenght is two times the width, what is the lenght and width?
Answer by rapaljer(4671)

Let x= width
2x= length

Area = W*L
Area = x*2x=50 Square Inches

2x^2=50

Divide both sides by 2:
x^2=25

There are two solutions: x=5 and x=-5. However, you have to reject any negative values, since you can't have a negative width.

x=5 inches = WIDTH
2x=2*5 = 10 inches = LENGTH

Check: Area = W*L=5*10= 50 square inches.

Merry Christmas!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus